OpenStudy (anonymous):

A person walks first at a constant speed of 5.10 m/s along a straight line from point circle a to point circle b and then back along the line from circle b to circle a at a constant speed of 2.95 m/s. a) what is the average speed over the entire trip ? b) What is the average velocity over the entire trip?

OpenStudy (anonymous):

a. Average Speed= (5.10+2.95)/2=4.025 m/s b. Average Velocity=(5.10-2.95)/2=1.075 m/s A -> B positive direction